Problem 2.59 You are driving eastbound on the interstate at 70 m/You observe that you are approaching a truck in your lane at a relative speed of 10 mi/h . Part A How fast is the truck moving relative to the highway? Express your answer in miles per hour to two significant figures. 60 mi/h Submit My Answers Give Up Correct Part B If the truck were instead traveling at this speed in the westbound lane, what would be the relative velocity between you and the truck? Express your answer in miles per hour to three significant figures. 60.0 mi/h SubmitM My Answers Give Up Incorrect; Try Again; 3 attempts remaining Continue

please answer part B.
incorrect answers : 80.0 and 60.0
Answer #1

Correct answer is 130 mi/hr

Explanation : where both are moving on eastbound lane

Vmy = 70 mi/hr , Vmy - Vtruck = 10 mi/hr

So Vtruck = 70 -10 = 60 mi/hr (taking eastbound as positive direction )

If truck is moving on westbound lane then Vtruck = -60mi/hr

Vrelative = Vmine - Vtruck = 70 -(-60) = 130mi/hr